You have purchased 15 shares of company DIY for $52 with 65% initial margin. The interest on the margin loan is 6%. How much is your return on your investment, if 3 months later DIY's price is $45? (Provide your answer in percentage rounded to two decimals, omitting the % sign)

Show work please. Will thumbs up if correct
Answer #1

Current Price = $ 52, Number of Shares Bought = 15, Total Worth of Purchase = 15 x 52 = $ 780

Initial Margin = 65 % of Purchase Value = 0.65 x 780 = $ 507

Borrowing = Purchase Value - Initial Margin = 780 - 507 = $ 273

Interest On Margin = 6 % of Borrowing = 0.06 x 273 = $ 16.38

Price of Shares after 3 months = $ 45

Portfolio Value after Three Months = 45 x 15 = $ 675

Return Value on Investment = 675 - 507 - 16.38 - 273 = - $ 121.38

Return on Investment = (-121.38/507) x 100 = - 23.94 %